This year Bobcat Company reports current E&P of negative $300,000 that accrued evenly throughout the year. At the beginning of the year, Bobcat’s accumulated E&P was $200,000. Bobcat distributed $200,000 to its sole shareholder, Melanie Rushmore, on June 30 of this year. Melanie’s tax basis in her Bobcat stock was $75,000. What is Bobcat’s balance in accumulated E&P on the first day of next year?
Answer #1

Melanie Rushmore reports a dividend of $ 51233.

He will first compute accumulated E&P as of June 30, 2011. The deficit in current E&P is allocated pro rata on a daily basis.

The deficit in current E&P = 300000 * ( 181 / 365 ) = 148767

accumulated E&P as of June 30, 2011 = 200000 - 148767 = $ 51233

Now, Bobcat’s balance in accumulated E&P on the first day of next year :-

accumulated E&P beginning 200000 - current E&P 300000 - Dividend paid 51233 =

accumulated E&P end of the year 151233

the deficit equals the deficit in current E&P that arose after the dividend payment =

300000 * ( 184 / 365 ) = 151233